### v5

# Thick isotopy property and the mapping class groups of Heegaard splittings

Abstract.
We give a necessary and sufficient condition for the fundamental group of the space of Heegaard splittings of an irreducible $3$-manifold to be finitely generated. The condition is exactly the conclusion of the thick isotopy lemma proved by Colding, Gabai and Ketover, which says that any isotopy of a Heegaard surface is achieved by a $1$-parameter family of surfaces with area bounded above by a universal constant and with some “thickness property”. We also prove that a Heegaard splitting of a hyperbolic or spherical $3$-manifold satisfies the condition if it is topologically minimal (in the sense of Bachman) and its disk complex has finitely generated homotopy group. In conclusion, such a Heegaard splitting has finitely generated mapping class group.
Key words and phrases:
3-manifold, Heegaard splitting, mapping class group, topologically minimal surface
1991 Mathematics Subject Classification:
57M60, 57M50
1. Introduction
Let $M$ be a closed orientable $3$-manifold. A Heegaard splitting is a decomposition of $M$ into two handlebodies along a closed embedded surface $\Sigma$. We will denote such a splitting of $M$ by $(M,\Sigma)$. In [26], Johnson and McCullough defined the space $\mathcal{H}(M,\Sigma)$ of Heegaard splittings equivalent to $(M,\Sigma)$ by $\mathrm{Diff}(M)/\mathrm{Diff}(M,\Sigma)$, where $\mathrm{Diff}(M)$ is the space of self-diffeomorphisms of $M$ and $\mathrm{Diff}(M,\Sigma)$ is its subspace consisting of maps that send $\Sigma$ to itself. For example, computing the 00-th homotopy group of $\mathcal{H}(M,\Sigma)$ is the same as classifying Heegaard splittings up to isotopy. Throughout the paper, we will focus only on the case that $M$ is irreducible. In [26], the $k$-th homotopy group of $\mathcal{H}(M,\Sigma)$ was computed for $k\geq 2$. On the other hand, $\pi_{1}(\mathcal{H}(M,\Sigma))$ is closely related to the mapping class group of a Heegaard splitting or the Goeritz group, and these groups are still mysterious. In this paper, we give a necessary and sufficient condition for $\pi_{1}(\mathcal{H}(M,\Sigma))$ to be finitely generated.
In [16], Colding, Gabai and Ketover found an effective algorithm to construct the complete list of Heegaard splittings of a non-Haken hyperbolic $3$-manifold. A key of their argument is the thick isotopy lemma ([16, Lemma 2.10]), which allows us to turn the computation of the 00-th homotopy group of the space of Heegaard splittings into a purely combinatorial problem, involving the (crudely) almost normal surface theory. The same strategy is also useful in computing $\pi_{1}(\mathcal{H}(M,\Sigma))$ as stated below. From now on, we fix a Riemannian metric on $M$. Let $\delta>0$. A surface $S$ in $M$ is said to be $\delta$-compressible if there exists a compressing disk $D$ for $S$ such that $\mathrm{diam}\,\partial D\leq\delta$. Otherwise $S$ is said to be $\delta$-locally incompressible.
Definition.
We say $(M,\Sigma)$ satisfies the thick isotopy property if the following holds. There exist $C>0$ and $\delta>0$, depending only on $\Sigma$ and the metric of $M$, such that any isotopy ${\{\Sigma_{t}\}}_{t\in I}$ with $\Sigma_{0}=\Sigma_{1}=\Sigma$ can be deformed within its homotopy class (as a loop in $\mathcal{H}(M,\Sigma)$) so that afterward for all $t\in I$,
  • $\mathrm{Area}(\Sigma_{t})<C$, and
  • $\Sigma_{t}$ is $\delta$-locally incompressible.
Theorem 1.1.
The fundamental group of $\mathcal{H}(M,\Sigma)$ is finitely generated if and only if $(M,\Sigma)$ satisfies the thick isotopy property.
Our second aim is to investigate what kind of Heegaard splitting satisfies the thick isotopy property. Let $S$ be a closed embedded surface in $M$ of genus at least $2$. The disk complex $\Gamma(S)$ of $S$ is defined to be the simplicial complex whose vertices are the isotopy classes of compressing disks for $S$, and whose $i$-simplices are $(i+1)$-tuples of vertices that admit disjoint representatives. In [3], Bachman introduced the concept of a topologically minimal surface as a generalization of several important classes of surfaces in a $3$-manifold, including incompressible surfaces and strongly irreducible surfaces.
Definition (Bachman [3]).
We say $S$ is topologically minimal if $\Gamma(S)=\emptyset$ or $\pi_{d-1}(\Gamma(S))\neq 1$ for some $d\in\mathbb{N}$. If $S$ is topologically minimal, the topological index of $S$ is defined to be the smallest number $d$ such that $\pi_{d-1}(\Gamma(S))\neq 1$.
Theorem 1.2.
Let $M$ be a hyperbolic or spherical $3$-manifold but not $S^{3}$. Let $(M,\Sigma)$ be a Heegaard splitting of $M$. Suppose that $\Sigma$ is a topologically minimal surface of index $d$. Furthermore, suppose that $\pi_{d-1}(\Gamma(\Sigma))$ is finitely generated if $d>1$. Then $(M,\Sigma)$ satisfies the thick isotopy property.
Here are a few remarks on the theorem. First, we note that the index of $\Sigma$ is $1$ if and only if $\Sigma$ is strongly irreducible, and so there are many examples of Heegaard splittings satisfying the assumption of the theorem. Unfortunately, we do not know if such examples exist when $d>1$. (This question can also be seen as the special case of Question 5.10 in [3].) However, the examples by Campisi-Rathbun [6] possibly satisfy the assumption. Building on the idea of Bachman-Johnson [4], they constructed examples of hyperbolic $3$-manifolds that contain a Heegaard surface with index $d$ for every $d>0$. Indeed, they proved that there exists a retraction from the disk complex to a sphere $P\subset\Gamma(\Sigma)$ of appropriate dimension. It is likely that we can arrange the construction so that such a sphere is in fact a deformation retract.
We also note that the assumption on $\pi_{d-1}(\Gamma(\Sigma))$ in Theorem 1.2 is used only in the proof of Lemma 5.3, and hence it can be replaced with any condition that implies Lemma 5.3. In particular, it would be interesting to search for examples of Heegaard splittings for which Lemma 5.3 holds without the assumption. In Section 6, we see that this is the case for infinitely many examples of Heegaard splittings of $(\mathrm{surface})\times I$. As a consequence, those Heegaard splittings satisfy the thick isotopy property.
The above theorems have an application to the theory of the mapping class group of a Heegaard splitting. For a Heegaard splitting $(M,\Sigma)$, its mapping class group $\mathrm{MCG}(M,\Sigma)$ is defined to be $\pi_{0}(\mathrm{Diff}(M,\Sigma))$.
Corollary 1.3.
If $M$ and $\Sigma$ are as in Theorem 1.2, then $\mathrm{MCG}(M,\Sigma)$ is finitely generated.
In Section 6, we also establish finite generation of the mapping class groups for infinitely many examples of Heegaard splittings of $(\mathrm{surface})\times I$.
There have been many efforts to find a finite generating set for the mapping class group of a Heegaard splitting. Possibly the most interesting is $\mathrm{MCG}(S^{3},\Sigma_{g})$, where $(S^{3},\Sigma_{g})$ is a standard genus $g$ Heegaard splitting of the $3$-sphere. It is known that $\mathrm{MCG}(S^{3},\Sigma_{g})$ is finitely generated for $g=2$ by [21] (see also [30]), and for $g=3$ by [20]. However, it is not known if the same is true for $g\geq 4$. On the other hand, a genus $\geq 2$ Heegaard surface in $S^{3}$ is topologically minimal by [2] or [7]. (But the disk complex is not of finite type. See Appendix A.) So there might be a good chance to improve our proof to remove the assumption that $M\neq S^{3}$. In fact, much of our argument is still valid when $M=S^{3}$: Lemma 5.5 below is the only place where the assumption $M\neq S^{3}$ is used essentially.
As another example, any genus $2$ weakly reducible Heegaard splitting has finitely presented mapping class group by [9, 10, 11, 12, 13, 1, 8]. A finite generating set for a genus $3$ Heegaard splitting of the $3$-torus is also known by [25]. While little has been known about the mapping class group of a Heegaard splitting of genus greater than $3$, an advantage of our approach is that it is applicable to arbitrarily high genus Heegaard splittings.
Organization of the paper
Section 2 is a preliminary towards the proof of Theorem 1.1, including the definition of a crudely almost normal surface. Theorem 1.1 is proved in Section 3. Section 4 is a quick introduction to min-max theory. In Section 5, we prove Theorem 1.2 and Corollary 1.3. In Section 6, we prove that infinitely many examples of Heegaard splittings of $({\rm surface})\times I$ satisfy the thick isotopy property, and as a consequence their mapping class groups are finitely generated. In Appendix A, we see that the disk complex of a Heegaard surface of $S^{3}$ is not homotopy equivalent to a finite simplicial complex.
Acknowledgements
The author would like to thank Professor Yuya Koda for his valuable comments, advice, and helpful conversations.
2. Normal surface theory
Throughout the paper, we will use the following notations:
  • $I:=[0,1]$.
  • For $r>0$, $B^{d}_{r}:=\{x\in\mathbb{R}^{d}\mid|x|\leq r\}$.
  • If $\mathcal{K}$ is a simplicial complex, we will denote by $\mathcal{K}^{i}$ its $i$-skeleton.
In this section, we recall some definitions and lemmas from [16]. Let $M$ be a closed orientable $3$-manifold. Let $\mathcal{T}$ be a triangulation of $M$.
Definition.
A closed embedded surface $S\subset M$ is crudely almost normal (with respect to $\mathcal{T}$) if the following are satisfied:
  1. (1)
    $S$ is transverse to any simplex of $\mathcal{T}$.
  2. (2)
    If $\tau$ is a $2$-simplex of $\mathcal{T}$, $S\cap\tau$ consists of finitely many arcs (with no circle component).
  3. (3)
    If $\sigma$ is a $3$-simplex of $\mathcal{T}$, $S\cap\sigma$ consists of finitely many disks but possibly with one exception: there may be exactly one $3$-simplex that contains exactly one unknotted annulus component.
A crudely almost normal surface will be called a crudely normal surface if it has no exceptional annulus component.
The weight of $S$ is defined to be $|S\cap\mathcal{T}^{1}|$. Let $S^{\prime}$ be another crudely almost normal surface. Then, $S$ and $S^{\prime}$ are said to be normally isotopic if they are isotopic through surfaces transverse to each simplex. We say $S^{\prime}$ is obtained from $S$ by a pinch if it is obtained from $S$ and a $2$-sphere in $M$ by connecting them with a tube. Such a move or its inverse will be called a pinch. Note that if $M$ is irreducible, a pinch can be achieved by an isotopy that contracts the $2$-sphere across a $3$-ball in $M$.
Lemma 2.1 ([16, Lemma 3.4]).
There are only finitely many normal isotopy classes of crudely almost normal surfaces with weight at most $L$.
A generic $\mathcal{T}$-isotopy is an isotopy $\{S_{t}\}_{t\in I}$ such that $S_{t}$ is transverse to $\mathcal{T}$ for all $t$ but finitely many points $10<t_{1},\ldots,t_{l}<$ 1. In addition, $S_{{t_{i}}+\epsilon}$ and $S_{{t_{i}}-\epsilon}$ differ by one of the moves shown in Figure 1: When $t$ moves from $t_{i}-\epsilon$ to $t_{i}+\epsilon$, $\Sigma_{t}$ passes either (0) a 00-simplex, (1) a $1$-simplex, (2a) a center tangency with a $2$-simplex or (2b) a saddle tangency. Furthermore, if $|S_{t}\cap\mathcal{T}^{1}|\leq L$ for $t\in I$, $\{S_{t}\}_{t\in I}$ is called a generic $L$-$\mathcal{T}$-isotopy. Finally, for a Riemannian $3$-manifold $M$, an isotopy $\{S_{t}\}_{t\in I}$ is said to be a $C$-isotopy if $\mathrm{Area}(S_{t})\leq C$ for $t\in I$.
[…]
Figure 1. When $t$ moves from $t_{i}-\epsilon$ to $t_{i}+\epsilon$, $\Sigma_{t}$ passes either (0) a 00-simplex, (1) a $1$-simplex, (2a) a center tangency with a $2$-simplex or (2b) a saddle tangency.
Lemma 2.2 ([16, Lemma 3.2]).
Let $\mathcal{T}$ be a triangulation of the Riemannian $3$-manifold $M$ with metric $\rho$, $L>0$ and $\varepsilon>0$. Then there exists $K(\mathcal{T},L,\epsilon,\rho)>0$ such that if $S$ is a closed embedded surface with $\mathrm{Area}(S)<C$, then $S$ is isotopic to a surface $S^{\prime}$ such that $|S^{\prime}\cap\mathcal{T}^{1}|<KC$ and the diameter of the trace of any point of the isotopy is at most $\varepsilon$.
If $F:S\times[0,1]\rightarrow M$ is a $C$-isotopy between surfaces $S_{0}$ and $S_{1}$ that are transverse to $\mathcal{T}$ of weight at most $L$, then there exists a generic $K(C+1)$-$\mathcal{T}$-isotopy $G$ from $S_{0}$ to $S_{1}$ such that, for all $x\in S$ and $t\in[0,1]$, $d(G(x,t),F(x,t))<\varepsilon$.
3. The proof of Theorem 1.1
In this section, we prove Theorem 1.1. It is not hard to see the necessity, i.e. forward implication, in the theorem. First, recall that a path in $\mathcal{H}(M,\Sigma)$ can be identified with an isotopy of a Heegaard surface. More precisely, as $\mathrm{Diff}(M)\to\mathrm{Diff}(M)/\mathrm{Diff}(M,\Sigma)=\mathcal{H}(M,\Sigma)$ is a fibration [26], any path $\alpha:I\to\mathcal{H}(M,\Sigma)$ lifts to $\tilde{\alpha}:I\to\mathrm{Diff}(M)$ and we can define an isotopy of a Heegaard surface by $\Sigma_{t}:=\tilde{\alpha}(t)(\Sigma)$. Conversely, if an isotopy of a Heegaard surface is given, it defines a path in $\mathcal{H}(M,\Sigma)$ via the isotopy extension theorem. Now if $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated, we can find a finite collection of isotopies of $\Sigma$ such that any isotopy representing an element of $\pi_{1}(\mathcal{H}(M,\Sigma))$ can be expressed as the product of isotopies in the collection. Thus, $(M,\Sigma)$ satisfies the thick isotopy property.
In the following, we prove the sufficiency of the theorem. Let $C>0$ and $\delta>0$ be the constants given in the definition of the thick isotopy property. Fix a triangulation $\mathcal{T}$ of $M$ such that
  • $\Sigma$ is crudely normal with respect to $\mathcal{T}$, and
  • any simplex of $\mathcal{T}$ has the diameter at most $\delta$.
Let $\{\Sigma_{t}\}_{t\in I}$ be any isotopy with $\Sigma_{0}=\Sigma_{1}=\Sigma$. By the argument in [16], we can convert $\Sigma_{t}$ to a crudely almost normal surface with respect to $\mathcal{T}$. Here is a very rough sketch of the argument. By assumption, $\mathrm{Area}(\Sigma_{t})\leq C$ and $\Sigma_{t}$ is $\delta$-locally incompressible for $t\in I$. By Lemma 2.2, $\Sigma_{t}$ is transverse to every simplex of $\mathcal{T}$ for all but finitely many points and $\Sigma_{t}$ has weight at most $L:=K(C+1)$. Using the $\delta$-locally incompressibility condition, for every $3$-simplex $\sigma$ of $\mathcal{T}$, we can pinch off and remove non-disk components of $\Sigma_{t}\cap\sigma$. (We note that a subtle situation may occur when $\Sigma_{t}$ passes through a tangency of type (2b). Around such a tangency, we may be forced to allow an unknotted annulus component. See [16, Proof of Lemma 3.6] for more details.) In summary, we have
Claim 1.
$\{\Sigma_{t}\}_{t\in I}$ can be deformed within its homotopy class (as the loop in the space $\mathcal{H}(M,\Sigma)$) so that afterward for all $t$ but finitely many points in $I$, $\Sigma_{t}$ is a crudely almost normal surface with weight bounded above by a universal constant $L>0$.
Proof.
This follows from [16, Lemma 3.6]. ∎
Consider the graph $\mathcal{G}$ such that each vertex of $\mathcal{G}$ corresponds to a normal isotopy class of crudely almost normal surfaces w.r.t. $\mathcal{T}$ with weight at most $L$, and each edge corresponds to one of the moves (0) – (2b) shown in Figure 1 or a pinch. By Lemma 2.1, $\mathcal{G}$ is a finite graph. In particular, $\pi_{1}(\mathcal{G})$ is finitely generated. By Claim 1, the natural homomorphism $\pi_{1}(\mathcal{G})\to\pi_{1}(\mathcal{H}(M,\Sigma))$ is a surjection. Thus, we conclude that $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated. ∎
4. The min-max theorem
This section is a quick introduction to the min-max theory of Simon-Smith [31], which will be used in the next section. One can consult e.g. [15, 17] for more details on this subject.
Let $M$ be a closed, orientable, Riemannian $3$-manifold. We will denote by $\mathscr{H}^{2}(\cdot)$ the $2$-dimensional Hausdorff measure on $M$.
Definition.
Let $X^{k}$ be a manifold. A family $\{\mathrm{\Sigma}_{t}\}_{t\in X}$ of closed subsets of $M$ is called a (genus $g$) sweep-out if it satisfies the following conditions:
  • $\mathrm{\Sigma}_{t}$ converges to $\mathrm{\Sigma}_{t_{0}}$ in the Hausdorff topology when $t\to t_{0}$.
  • $\mathscr{H}^{2}(\mathrm{\Sigma}_{t})\rightarrow\mathscr{H}^{2}(\mathrm{\Sigma}_{t_{0}})$ when $t\rightarrow t_{0}$.
  • $\Sigma_{t}$ is a closed genus $g$ surface in $M$ if $t\in\mathrm{int}\,X$. On the other hand, if $t\in\partial X$, $\Sigma_{t}$ is a closed surface of $\mbox{genus}\leq g$ plus finitely many arcs.
  • $\mathrm{\Sigma}_{t}$ varies smoothly for $t\in\mathrm{int}\,X$.
For later use, we restrict ourselves to the case that $X=I\times B^{d}$. Consider the subspace $\mathscr{I}$ of $C^{\infty}(M\times(I\times B^{d}),M)$ consisting of those maps $\psi$ such that
  1. ($\mathrm{i}$)
    $\psi(\cdot,t)$ is a diffeomorphism of $M$ for $t\in I\times B^{d}$, and
  2. ($\mathrm{ii}$)
    $\psi(\cdot,t)=\mathrm{id}_{M}$ for $t\in\partial I\times B^{d}$.
Let $\mathscr{I}_{0}\subset\mathscr{I}$ be the component containing the map $\psi_{0}$ given by $\psi_{0}(x,t):=x$. Given a sweep-out $\{\mathrm{\Sigma}_{t}\}_{t\in I\times B^{d}}$, define the collection $\Pi_{\{\mathrm{\Sigma}_{t}\}}$ of sweep-outs by
$\Pi_{\{\mathrm{\Sigma}_{t}\}}:=\left\{\psi(\Sigma_{t},t)\}_{t\in I\times B^{d}}\mid\psi\in\mathscr{I}_{0}\right\}.$
The width of $\Pi_{\{\mathrm{\Sigma}_{t}\}}$ is defined by
$W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M):=\inf_{\{\Lambda_{t}\}\in\Pi_{\{\mathrm{\Sigma}_{t}\}}}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\Lambda_{t}).$
A sequence $\{\mathrm{\Sigma}_{t}^{i}\}_{t\in X}$ ($i\in\mathbb{N}$) of sweep-outs in $\Pi_{\{\mathrm{\Sigma}_{t}\}}$ is a minimizing sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\sup_{t\in I\times B^{d}}\mathscr{H}^{2}(\mathrm{\Sigma}_{t}^{i})$. Furthermore, a sequence $\{\mathrm{\Sigma}_{t_{i}}^{i}\}_{i\in\mathbb{N}}$ is a min-max sequence if $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\lim_{i\to\infty}\mathscr{H}^{2}(\mathrm{\Sigma}_{t_{i}}^{i})$.
Simon-Smith’s min-max theorem is the following theorem. (The following statement can be found in [16] with minor modification, see [15, 18, 27] for the proof and also [16, Appendix] for the multi-parameter case.)
Theorem 4.1 (cf. [16, Theorem 2.1]).
Given a sweep-out $\{\mathrm{\Sigma}_{t}\}_{t\in I\times B^{d}}$ of genus-$g$ surfaces, if
(4.1) $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)>\sup_{t\in\partial I\times B^{d}}\mathscr{H}^{2}(\mathrm{\Sigma}_{t}),$
then there exists a min-max sequence $\mathrm{\Sigma}_{i}:=\mathrm{\Sigma}_{t_{i}}^{i}$ such that
[…]
where $\Gamma_{i}$ are smooth closed embedded minimal surfaces and $n_{i}$ are positive integers. Moreover, after performing finitely many compressions on $\mathrm{\Sigma}_{i}$ and discarding some components, each connected component is isotopic to one of the $\Gamma_{i}$ or to a double cover of one of the $\Gamma_{i}$. We have the following genus bounds with multiplicity:
(4.3) $\sum_{i\in\mathscr{O}}n_{i}g(\Gamma_{i})+\frac{1}{2}\sum_{i\in\mathscr{N}}n_{i}(g(\Gamma_{i})-1)\leq g,$
where $\mathscr{O}$ denotes the subcollection of $\Gamma_{i}$ that is orientable and $\mathscr{N}$ denotes those $\Gamma_{i}$ that are nonorientable, and where $g(\Gamma_{i})$ denotes the genus of $\Gamma_{i}$ if it is orientable, and the number of cross-caps that one attaches to a sphere to obtain a homeomorphic surface if $\Gamma_{i}$ is nonorientable.
Lemma 4.2.
Let $M$ be a hyperbolic or spherical $3$-manifold. If $\{\Sigma_{t}\}_{t\in I\times B^{d}}$ is a genus $g$ sweep-out satisfying $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)>\sup_{t\in\partial I\times B}\mathscr{H}^{2}(\mathrm{\Sigma}_{t})$, then $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)\leq 8\pi(g+1)$.
Proof.
If $M$ is hyperbolic, $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)\leq 4\pi(g-1)$ by [17, Lemma 9.4]. So, we prove the lemma when $M$ is spherical. Theorem 4.1 shows that $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\sum_{i=1}^{k}n_{i}\mathrm{Area}(\Gamma_{i})$ for some embedded minimal surfaces $\Gamma_{i}$ ($1\leq i\leq k$). By Frankel’s theorem [19], the min-max limit is in fact connected and we can express the width as $W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=n\mathrm{Area}(\Gamma)$. If $\Gamma$ is non-orientable, its double cover is stable by [32, Theorem 7.2]. But this is impossible because $S^{3}$ with the standard metric and thus its quotient cannot contain a stable minimal surface. So $\Gamma$ must be orientable and again by [32, Theorem 7.2], the multiplicity $n$ must be one. This together with Choi-Schoen’s area bound [14] for a minimal surface in a spherical $3$-manifold implies
$W(\Pi_{\{\mathrm{\Sigma}_{t}\}},M)=\mathrm{Area}(\Gamma)\leq 8\pi\left(\frac{2}{|\pi_{1}(M)|}-\frac{\chi(\Gamma)}{2}\right)\leq 8\pi(g+1).$
5. Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Suppose that $M$ is a hyperbolic or spherical $3$-manifold but not $S^{3}$, and $\Sigma$ is a genus $g$ Heegaard surface of topological index $d>0$. Furthermore, if $d>1$, we assume that $\pi_{d-1}(\Gamma(\Sigma))$ is finitely generated.
Step 1: The definitions of $C$ and $\delta$
Let $\{\varphi_{k}:S^{d-1}\to\Gamma(\Sigma)\mid k=1,\ldots,n\}$ be a collection of maps that represents a finite generating set of $\pi_{d-1}(\Gamma(\Sigma))$. Put $B=B^{d}_{1}$. For each $\varphi_{k}$, we consider the sweep-out $\{\Sigma^{k}_{s}\}$ parametrized by $B$ and given below. (The same construction can be found in [3].)
Take a triangulation $\mathcal{K}$ of $S^{d-1}$ for which $\varphi_{k}$ is simplicial. For each vertex $s$ of $\mathcal{K}$, choose a representative $D_{s}$ of $\varphi_{k}(s)\in\Gamma(\Sigma)$ so that if $s$ and $s^{\prime}$ are in the same simplex of $\mathcal{K}$, $D_{s}$ and $D_{s^{\prime}}$ are disjoint. Let $\widetilde{\mathcal{K}}$ be the cone over $\mathcal{K}$. So, $\widetilde{\mathcal{K}}$ is a triangulation of $B$. Roughly, $\{\Sigma^{k}_{s}\}_{s\in B}$ is defined as follows. For the center 00 of $B$, set $\Sigma^{k}_{0}=\Sigma$. For each vertex $s$ of $\mathcal{K}$, let $\Sigma^{k}_{s}$ be the result of compressing $\Sigma$ across $D_{s}$. Furthermore, for a barycenter $s$ of an $l$-simplex of $\mathcal{K}$ that is spanned by vertices $s_{0},\ldots,s_{l}\in\mathcal{K}^{0}$, define $\Sigma^{k}_{s}$ to be the result of compressing $\Sigma$ simultaneously across $D_{s_{i}}$. For general $s\in B$, we define $\Sigma^{k}_{s}$ as a linear interpolation of the above construction.
Here is a more formal definition. Let $\tilde{\sigma}$ be a $d$-simplex of $\tilde{\mathcal{K}}$ and let $0=s_{0},s_{1},\ldots,s_{d}$ be the vertices of $\tilde{\sigma}$. Denote by $\sigma$ the $(d-1)$-simplex spanned by the vertices $s_{1},\ldots,s_{d}$. For $1\leq i\leq d$, fix a cylinder $N(D_{s_{i}})\cong B^{2}_{1}\times[-1,1]$ in $M$ such that $B^{2}_{1}\times 0=D_{s_{i}}$ and $N(D_{s_{i}})\cap\Sigma=\partial B^{2}_{1}\times[-1,1]$. Furthermore, $N(D_{s_{i}})$ can be chosen to be mutually disjoint. Let $\rho_{\tau}:B^{2}_{1}\times[-1,1]\to B^{2}_{1}\times[-1,1]$ be the map given by
$\rho_{\tau}(x,u)=((1-\tau b(u))x,u)$
for $(x,u)\in B^{2}_{1}\times[-1,1]$ and $\tau\in I$. Here, $b(u)$ is a bump function. Thus, $\rho_{\tau}$ shrinks the annulus $\partial B^{2}_{1}\times[-1+\epsilon,1-\epsilon]$ to an arc $0\times[-1+\epsilon,1-\epsilon]$. Fix an identification $\iota_{s_{i}}:B^{2}_{1}\times[-1,1]\to N(D_{s_{i}})$ for $1\leq i\leq d$, and define the homotopy $\eta^{s_{i}}_{\tau}:M\to M$ to be
$\eta^{s_{i}}_{\tau}:=\iota_{s_{i}}\circ\rho_{\tau}\circ\iota^{-1}_{s_{i}}$
on $N(D_{s_{i}})$ and to be the identity on the complement of a small neighborhood of $N(D_{s_{i}})$. Regard $\tilde{\sigma}$ as a $d$-cube in $\mathbb{R}^{d}$ so that $0=s_{0}$, the $i$th unit vector represents $s_{i}$, and the other corners correspond to the barycenters of the simplices contained in $\sigma$. Then, every point $s\in\tilde{\sigma}$ can be written as $s=\tau_{1}s_{1}+\cdots+\tau_{d}s_{d}$ for $\tau_{1},\ldots,\tau_{d}\in I$. Now we define $\Sigma^{k}_{s}$ as the image of $\Sigma$ by the composition
$\eta^{s_{1}}_{\tau_{1}}\circ\cdots\circ\eta^{s_{d}}_{\tau_{d}}.$
($\eta^{s_{i}}_{\tau_{i}}$’s are commutative with each other since their supports are mutually disjoint.) Since the above construction is consistent on every intersection between adjacent simplices, we obtain the sweep-out $\{\Sigma^{k}_{s}\}_{s\in B}$.
For simplicity, given a surface $T$ in $M$, we define
$\gamma(T):=\min\{\mathrm{diam}\,\partial D\mid\mbox{$D$ is a compressing disk for $T$}\}.$
We write $\mathrm{inj}(M)$ for the injectivity radius of $M$. Now define
$C:=\max\left\{\max_{s\in B}\mathscr{H}^{2}(\Sigma^{1}_{s}),\ldots,\max_{s\in B}\mathscr{H}^{2}(\Sigma^{n}_{s}),8\pi(g+1)\right\}+1,$
and
$\delta:=\min\{3^{-(1+2+\cdots+d+(d+1))}\cdot\mathrm{inj}(M),\gamma(\Sigma)/2\}.$
Step 2: Lemmas
Let $k\in\{1,\ldots,n\}$. Note that by the isotopy extension theorem, we can fix a diffeomorphism $f_{s}:\Sigma\to\Sigma^{k}_{s}$ for $s\in\mathrm{int}\,B$ simultaneously. In the next lemma, which follows from the definition of $\{\Sigma^{k}_{s}\}_{s\in B}$, we identify $\Gamma(\Sigma)$ with $\Gamma(\Sigma^{k}_{s})$ through these diffeomorphisms.
Lemma 5.1.
Let $S=\partial B^{d}_{1-\epsilon}$. If $\epsilon$ is small enough, $\Sigma^{k}_{s}$ is $\delta$-compressible for $s\in S$. Furthermore, if $\mathcal{S}$ is a triangulation of $S$ such that the diameter of each simplex is small enough, and if we define $\psi:\mathcal{S}^{0}\to\Gamma(\Sigma)^{0}$ by sending $s\in\mathcal{S}^{0}$ to one of $\delta$-compressing disks for $\Sigma_{s}$ (and applying $f_{s}^{-1}$), then $\psi$ determines the simplicial map $\mathcal{S}\to\Gamma(\Sigma)$ homotopic to $\varphi_{k}$.
Define $U_{0}\subset B$ to be the set of points $s$ such that $\Sigma^{k}_{s}$ is $\delta$-compressible.
Lemma 5.2.
$B\setminus U_{0}$ is a star-shaped region.
Proof.
The proof is by contradiction. Suppose that there exists a line segment $\ell\subset B$ connecting 00 with $s\in B\setminus U_{0}$ such that $\ell$ contains a point $u\in U_{0}$. We can find a compressing disk $D$ for $\Sigma^{k}_{u}$ with $\mathrm{diam}\,\partial D\leq\delta$. Let $Q$ be a $3$-ball of diameter $\leq\delta$ that contains $\partial D$. After perturbing $Q$, $\partial Q$ intersects $\Sigma^{k}_{0}=\Sigma$, $\Sigma^{k}_{s}$ and $\Sigma^{k}_{u}$ transversely. If $Q\cap\Sigma_{s}$ contains circles that are essential in $\Sigma^{k}_{s}$, one of such circles bounds a compressing disk for $\Sigma^{k}_{s}$. As $\Sigma^{k}_{s}$ is $\delta$-locally incompressible, this case cannot occur. Thus, all the circles in $Q\cap\Sigma^{k}_{s}$ are inessential in $\Sigma^{k}_{s}$. Similarly, all the circles in $Q\cap\Sigma$ are inessential in $\Sigma$. By the innermost disk argument, we can isotope $Q$ so that $Q$ is contained in the region between $\Sigma$ and $\Sigma^{k}_{s}$, which is diffeomorphic to $\Sigma\times I$. Note that $Q$ still intersects $\Sigma^{k}_{u}$ so that one of the circles in $Q\cap\Sigma^{k}_{u}$ is essential in $\Sigma^{k}_{u}$. Thus, $\Sigma^{k}_{u}$ is compressible in the product region. This is a contradiction because $\Sigma^{k}_{u}$ is isotopic to a level surface in the product region. ∎
Step 3: Extending the sweep-out
Fix a nontrivial map $\varphi:S^{d-1}\to\Gamma(\Sigma)$ once. We may assume that $\varphi=\varphi_{1}$ and set $\{\Sigma_{0s}\}_{s\in B}:=\{\Sigma^{1}_{s}\}_{s\in B}$.
Let $h_{t}:\Sigma\to M$ be any isotopy with $h_{0}(\Sigma)=h_{1}(\Sigma)=\Sigma$. To prove Theorem 1.2, we must show that there exists an isotopy $h^{\prime}_{t}$ equivalent to $h_{t}$ such that for $t\in I$
  • $\mathrm{Area}(h^{\prime}_{t}(\Sigma))<C$, and
  • $h^{\prime}_{t}(\Sigma)$ is $\delta$-locally incompressible.
By the isotopy extension theorem, $h_{t}:\Sigma\to M$ extends to $\tilde{h}_{t}:M\to M$. Now define $\Sigma_{ts}$ for $t\in I$ and $s\in B$ by
$\Sigma_{ts}:=\tilde{h}_{t}(\Sigma_{0s}).$
Lemma 5.3.
$\{\Sigma_{ts}\}_{(t,s)\in I\times B}$ extends to $\{\Sigma_{ts}\}_{(t,s)\in[0,2]\times B}$ such that
  1. (1)
    $\Sigma_{t0}=\Sigma$ for $t\in[1,2]$,
  2. (2)
    $\mathrm{Area}(\Sigma_{2s})<C$ for $s\in B$,
  3. (3)
    $B\setminus U_{2}$ is a star-shaped region, where $U_{2}$ is the set of those points $s$ such that $\Sigma_{2s}$ is $\delta$-compressible.
Proof.
First, note that there is a natural action of $\mathrm{Diff}(M,\Sigma)$ on $\Gamma(\Sigma)$, which induces the action on $[S^{d-1},\Gamma(\Sigma)]$. By construction, $\{\Sigma_{1s}\}_{s\in B}$ is the image of $\{\Sigma_{0s}\}_{s\in B}$ by $\tilde{h}_{1}$. In other words, $\{\Sigma_{1s}\}_{s\in B}$ can be recovered from $\tilde{h}_{1}\cdot\varphi$ as follows. We repeat the same construction as in Step 1. Take a triangulation $\mathcal{K}$ of $S^{d-1}$ such that $\varphi:S^{d-1}\to\Gamma(\Sigma)$ is simplicial. For each vertex $s$ of $\mathcal{K}$, choose a compressing disk $D_{s}$ for $\Sigma$ that represents $\varphi(s)$. Then, $\tilde{h}_{1}(D_{s})$ is a compressing disk that represents $\tilde{h}_{1}\cdot\varphi(s)$. Fix an identification $\kappa_{s}:B^{2}_{1}\times[-1,1]\to\tilde{h}_{1}(N(D_{s}))$ for each $s\in\mathcal{K}^{0}$, and define the homotopy $\theta^{s}_{\tau}:M\to M$ to be
$\theta^{s}_{\tau}:=\kappa_{s}\circ\rho_{\tau}\circ\kappa_{s}^{-1}$
on $\tilde{h}_{1}(N(D_{s}))$ and to be the identity on the complement of a small neighborhood of $\tilde{h}_{1}(N(D_{s}))$. On each simplex $\tilde{\sigma}$ of $\widetilde{\mathcal{K}}$ with vertices $0=s_{0},s_{1},\ldots,s_{d}$, define $\Sigma_{1s}$ by the image of the homotopy
$\theta^{s_{1}}_{\tau_{1}}\circ\cdots\circ\theta^{s_{d}}_{\tau_{d}}:M\to M,$
where $s=\tau_{1}s_{1}+\cdots+\tau_{d}s_{d}$ and $\tau_{1},\ldots,\tau_{d}\in I$. Note that the above construction contains an ambiguity regarding the choice of an identification $B^{2}_{1}\times[-1,1]\to\tilde{h}_{1}(N(D_{s}))$: we have to see that $\kappa_{s}$ coincide with the image of the “standard” one, that is $\tilde{h}_{1}\circ\iota_{s}$. But these two maps are isotopic by the uniqueness of tubular neighborhoods, and hence the corresponding sweep-outs can be interpolated via this isotopy. In this way, we recover the sweep-out $\{\Sigma_{1s}\}_{s\in B}$.
By assumption, there is a homotopy $\Phi:[1,2]\times S^{d-1}\to\Gamma(\Sigma)$ such that $\Phi_{1}=\tilde{h}_{1}\cdot\varphi$ and $\Phi_{2}$ is a product of $\varphi_{k}$’s. By the relative simplicial approximation theorem (see e.g. [33]), we can extend $\{1\}\times\mathcal{K}$ to a triangulation $\mathcal{L}$ of $[1,2]\times S^{d-1}$ ($=[1,2]\times\partial B$) such that $\Phi$ is simplicial with respect to $\mathcal{L}$. In the following, we will construct a sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[1,2]\times B}$ that “shadows” the homotopy $\Phi$.
The construction is similar to that in Step 1. Extend the triangulation $\mathcal{L}$ over $\partial([1,2]\times B)$ by adding the two points $(1,0)$ and $(2,0)$ as vertices. Regarding $[1,2]\times B$ as a cone over $\partial([1,2]\times B)$, we obtain a triangulation $\widetilde{\mathcal{L}}$ of $[1,2]\times B$. For each vertex $v=(t,s)$ on $\mathcal{L}$, choose a representative $D_{v}$ of $\Phi(v)$ so that if $v$ and $v^{\prime}$ are in the same simplex, $D_{v}$ and $D_{v^{\prime}}$ are disjoint. For $v=(1,0)$ or $(2,0)$, let $D_{v}=\emptyset$.
As before, fix an identification $\kappa_{v}:B^{2}_{1}\times[-1,1]\to N(D_{v})$ for each vertex $v$ of $\mathcal{L}$, and define $\theta^{v}_{\tau}:M\to M$ to be $\kappa_{v}\circ\rho_{\tau}\circ\kappa^{-1}_{v}$ on $N(D_{v})$ and to be the identity outside a small neighborhood of $N(D_{v})$. It suffices to construct the sweep-out on every simplex $\tilde{\sigma}$ of $\widetilde{\mathcal{L}}$. Let $v_{0},\ldots,v_{d+1}$ be the vertices of $\tilde{\sigma}$ and regard $\tilde{\sigma}$ as a $(d+1)$-cube in Euclidean space with $v_{0}=0$. For $x=\tau_{1}v_{1}+\cdots+\tau_{d+1}v_{d+1}\in\tilde{\sigma}$, $\tau_{1},\ldots,\tau_{d+1}\in I$, define $\Sigma_{x}$ as the image of $\Sigma$ by the composition
$\theta^{v_{1}}_{\tau_{1}}\circ\cdots\circ\theta^{v_{d+1}}_{\tau_{d+1}}.$
Note that for $v\in\{2\}\times S^{d-1}$, the disk $D_{v}$ representing $v$ may not coincide with the standard one which has been chosen in Step 1. Of course, $D_{v}$ is isotopic to the standard choice, and in addition if $v_{1},\ldots,v_{l}$ be vertices adjacent to $v$ and all the $D_{v_{i}}$ have already been in the standard position, then we can isotope $D_{v}$ to the standard position in the complement of $D_{v_{i}}$. Thus, after these isotopies, we may assume that all the $D_{v}$ coincide with the standard one. By the uniqueness of tubular neighborhoods, we can make $N(D_{v})$ and $\theta^{v}_{\tau}$ standard as well.
This sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[0,2]\times B}$ satisfies the desired property. Indeed, if $\ell\subset B$ is a radial segment in $B$, $\{\Sigma_{2s}\}_{s\in\ell}$ appears in some $\{\Sigma^{k}_{s}\}_{s\in B}$ as a subfamily. Thus (2) and (3) hold. (1) is obvious from the construction. ∎
By Theorem 4.1, we have the following lemma.
Lemma 5.4.
$\{\Sigma_{ts}\}_{(t,s)\in[0,2]\times B}$ can be modified so that afterward $\mathscr{H}^{2}(\Sigma_{ts})<C$ for $t\in[0,2]$ and $s\in B$.
Step 4: Lifting a submanifold of $[0,2]\times B$ to $\Gamma(\Sigma)$
We fix some notation and terminology. Let us fix diffeomorphisms $f_{ts}:\Sigma_{00}\to\Sigma_{ts}$ for $(t,s)\in I\times\mathrm{int}\,B$ simultaneously via the isotopy extension theorem. We can identify $\Gamma(\Sigma_{ts})$ with $\Gamma(\Sigma_{00})$ ($=\Gamma(\Sigma)$) through this identification. We say $\Sigma_{ts}$ and $\Sigma_{t^{\prime}s^{\prime}}$ are $\epsilon$-close if $d_{M}(f_{ts}(x),f_{t^{\prime}s^{\prime}}(x))<\epsilon$ for any $x\in\Sigma_{00}$. Finally, define $U\subset[0,2]\times B$ to be the set of all points $(t,s)$ such that $\Sigma_{ts}$ is $\delta$-compressible.
Lemma 5.5.
Let $Y$ be a $d$-manifold embedded in $U$. Let $\epsilon<\delta$. Suppose that $\mathcal{Y}$ is a triangulation of $Y$ such that if $y,y^{\prime}$ are in the same simplex of $\mathcal{Y}$, then $\Sigma_{y}$ is $\epsilon$-close to $\Sigma_{y^{\prime}}$. Let $\mathcal{Y}^{\prime}$ be the barycentric subdivision of $\mathcal{Y}$. If we define the map $\psi:\mathcal{Y}^{0}\to\Gamma(\Sigma)^{0}$ by sending $y\in\mathcal{Y}^{0}$ to one of $\delta$-compressing disks for $\Sigma_{y}$ and applying $f_{y}^{-1}$, it extends to a simplicial map $\bar{\psi}:\mathcal{Y}^{\prime}\to\Gamma(\Sigma)$.
Proof.
To extend $\psi$ to a simplicial map $\bar{\psi}:\mathcal{Y}^{\prime}\to\Gamma(\Sigma_{00})$, it suffices to find a collection $\{D_{y}\mid y\in\mathcal{Y}^{\prime 0}\}$ of disks with the following property.
  • For $y\in\mathcal{Y}^{\prime 0}$, $D_{y}$ is a compressing disk for $\Sigma_{y}$.
  • If $y$ and $y^{\prime}$ are in the same simplex of $\mathcal{Y}^{\prime}$, then $f_{y}^{-1}(D_{y})$ and $f_{y^{\prime}}^{-1}(D_{y^{\prime}})$ are disjoint (i.e. they span a $1$-simplex in $\Gamma(\Sigma_{00})$).
Indeed, if such a collection of disks exists, we can define $\bar{\psi}$ by $\bar{\psi}(y)=f_{y}^{-1}(D_{y})$. The proof is by induction: we will show
Claim.
If $D_{y^{\prime}}$ has already been defined for $y^{\prime}$ the barycenter of any $(i-1)$-simplex of $\mathcal{Y}$ and $\mathrm{diam}\,\partial D_{y^{\prime}}<3^{1+\cdots+i}\cdot\delta$ holds, then we can find $D_{y}$ for $y$ the barycenter of any $i$-simplex $\mathcal{Y}$ such that $\mathrm{diam}\,\partial D_{y}<3^{1+\cdots+i+(i+1)}\cdot\delta$.
Let $\sigma$ be an $i$-simplex of $\mathcal{Y}$ and let $y$ be the barycenter of $\sigma$. Let $y_{1},\ldots,y_{2^{i+1}-1}=y$ be the vertices of $\mathcal{Y}^{\prime}$ that are contained in $\sigma$. By induction, for $21\leq j\leq 2^{i+1}-$ 2, $D_{y_{j}}$ has already been defined and the diameter of $\partial D_{y_{j}}$ is less than $3^{1+\cdots+i}\cdot\delta$. As $\Sigma_{y_{j}}$ and $\Sigma_{y}$ are $\epsilon$-close, the image of $\partial D_{y_{j}}$ on $\Sigma_{y}$ has diameter less than $3^{1+\cdots+i}\cdot\delta+2\epsilon$. In what follows, we work on a single surface, say $\Sigma_{y}$, rather than multiple surfaces. We will not distinguish between $D_{y_{j}}$ and its image on $\Sigma_{y}$ from their notation.
After relabeling $y_{j}$’s if necessary, we can assume that there exists a number $k$ ($21\leq k\leq 2^{i+1}-$ 2) satisfying the following: there exists a metric ball $Q$ with $\mathrm{diam}\,Q<k(3^{1+\cdots+i}\cdot\delta+2\epsilon)$ such that $\bigcup_{j=1}^{k}\partial D_{y_{j}}$ is contained in $Q$ while $\bigcup_{j=k+1}^{2^{i+1}-2}\partial D_{y_{j}}$ is in the complement of $Q$. Note that $Q$ is a genuine $3$-ball because
$\mathrm{diam}\,Q$ $<k(3^{1+\cdots+i}\cdot\delta+2\epsilon)$
$\leq(2^{i+1}-2)\cdot(3^{1+\cdots+i}\cdot\delta+2\delta)$
$<3^{1+\cdots+(i+1)}\cdot\delta$
$\leq\mathrm{inj}(M).$
After perturbing $Q$, we assume that $\partial Q$ intersects $\Sigma_{y}$ transversely. We can find a circle in $\partial Q\cap\Sigma_{y}$ that is essential in $\Sigma_{y}$. Indeed, if all the circles in $\partial Q\cap\Sigma_{y}$ were inessential, by the innermost disk argument, $\Sigma_{y}$ could be isotoped so that afterward $\Sigma_{y}\subset Q$. This is impossible because $\Sigma_{y}$ is a Heegaard surface and $M$ is not a $3$-sphere. So one of the circles in $\partial Q\cap\Sigma_{y}$ bounds a compressing disk for $\Sigma_{y}$. Define $D_{y}$ as such a disk. By definition, $D_{y}\cap D_{y_{j}}=\emptyset$ for $21\leq j\leq 2^{i+1}-$ 2 and $\mathrm{diam}\,\partial D_{y}<3^{1+\cdots+(i+1)}\cdot\delta$, which proves the claim. ∎
Step 5: The conclusion
We now finish the proof of Theorem 1.2. If $(0,0)$ and $(2,0)$ can be connected by a path in $[0,2]\times B$ without meeting $U$, it defines an isotopy $h^{\prime}_{t}$ with the desired property, proving the theorem. Thus, it suffices to show that $[0,2]\times B\setminus U$ is path-connected. We will prove this by contradiction. Recall that $S=\partial B^{d}_{1-\epsilon}$.
Claim 2.
There exists a compact orientable $d$-manifold $Y$ in $U$ with $\partial Y=0\times S$.
Proof.
Consider the map $f:[0,2]\times\mathrm{int}\,B\to\mathbb{R}$ given by $f(t,s):=\gamma(\Sigma_{ts})$. Since $\Sigma_{ts}$ varies smoothly for $(t,s)\in[0,2]\times\mathrm{int}\,B$, $f$ is a continuous function. By the smooth approximation theorem, $f$ is approximated by a smooth map $f^{\prime}$. Let $r\in\mathbb{R}$ be a regular value of $f^{\prime}$ just below $\delta$. By assumption, one of the components of $f^{\prime-1}(r)$, say $Y^{\prime}$, separates $(0,0)$ from $(2,0)$. On the other hand, by construction, if $(t,s)$ is close enough to $[0,2]\times\partial B$, then $f^{\prime}(t,s)<r$. This implies that $\partial Y^{\prime}\subset\{0,2\}\times B$. By Lemmas 5.2 and 5.3 (3), $Y^{\prime}$ extends to a $d$-manifold $Y$ in $U$ with $\partial Y=0\times S$. ∎
Pick a triangulation $\mathcal{Y}$ of $Y$ such that the diameter of any simplex of $\mathcal{Y}$ is small enough. Let $\mathcal{Y}^{\prime}$ be the barycentric subdivision of $\mathcal{Y}$. By Lemma 5.5, we can find a simplicial map $\bar{\psi}:\mathcal{Y}^{\prime}\to\Gamma(\Sigma)$. By Lemma 5.1, the restriction of $\bar{\psi}$ on $\partial Y$ must be homotopic to $\varphi$. Thus, $\varphi$ is homologically trivial. If $d\neq 2$, the Hurewicz theorem implies that $\varphi$ is homotopically trivial, contradicting the choice of $\varphi$.
If $d=2$, we can deduce a contradiction as follows. Let $V$ and $W$ be the handlebodies in $M$ bounded by $\Sigma$. Denote by $\Gamma_{V}(\Sigma)$ (resp. $\Gamma_{W}(\Sigma)$) the subcomplex of $\Gamma(\Sigma)$ spanned by compressing disks for $\Sigma$ that lie in $V$ (resp. $W$). Furthermore, denote by $\Gamma_{VW}(\Sigma)$ the union of all simplices that contain vertices in both $\Gamma_{V}(\Sigma)$ and $\Gamma_{W}(\Sigma)$. Thus, $\Gamma(\Sigma)=\Gamma_{V}(\Sigma)\cup\Gamma_{VW}(\Sigma)\cup\Gamma_{W}(\Sigma)$. Recall that the choice of $\varphi$ is arbitrary as long as it is homotopically nontrivial. By Claim 2.7 in [3], we can assume that $\varphi$ is represented by a loop $\gamma$ in $\Gamma(\Sigma)$ with the following properties:
  1. (a)
    $\gamma$ can be expressed as $e\cup\gamma_{V}\cup e^{\prime}\cup\gamma_{W}$, where $e$, $e^{\prime}$ are edges in $\Gamma_{VW}(\Sigma)$ while $\gamma_{V}$, $\gamma_{W}$ are paths in $\Gamma_{V}(\Sigma)$ and $\Gamma_{W}(\Sigma)$, respectively.
  2. (b)
    $e$ is in the different component of $\Gamma_{VW}(\Sigma)$ from $e^{\prime}$.
By definition, $\bar{\psi}(\partial Y)=\gamma$. Note that $\bar{\psi}^{-1}(\Gamma_{V}(\Sigma))\cap\bar{\psi}^{-1}(\Gamma_{W}(\Sigma))=\emptyset$. This along with (a) implies that there exists an arc in $\bar{\psi}^{-1}(\Gamma_{VW}(\Sigma))$ connecting $\bar{\psi}^{-1}(e)$ and $\bar{\psi}^{-1}(e^{\prime})$. This contradicts (b) and completes the proof of Theorem 1.2. ∎
Proof of Corollary 1.3
We can now prove Corollary 1.3. By Theorems 1.1 and 1.2, $\pi_{1}(\mathcal{H}(M,\Sigma))$ is finitely generated, and this group projects onto $\mathrm{Isot}(M,\Sigma)$, the subgroup of $\mathrm{MCG}(M,\Sigma)$ that consists of maps $(M,\Sigma)\to(M,\Sigma)$ isotopic to $\mathrm{id}_{M}$. Since $\mathrm{MCG}(M)$ is a finite group, $\mathrm{Isot}(M,\Sigma)$ has finite index in $\mathrm{MCG}(M,\Sigma)$. Thus, $\mathrm{MCG}(M,\Sigma)$ is also finitely generated.
6. Examples: Heegaard splittings of $({\rm surface})\times I$
Let $F$ be a closed orientable surface of genus $g\geq 2$, and set $P:=F\times[-1,1]$. For $n\in\mathbb{N}$, we consider the Heegaard surface $\Sigma^{n}=\Sigma$ for $P$ constructed as follows. Let $-1<r_{1}<\cdots<r_{n+1}<1$ and $F_{i}:=F\times\{r_{i}\}$. Fix a vertical arc $a_{i}$ in $F\times[r_{i},r_{i+1}]$ connecting $F_{i}$ to $F_{i+1}$ and let $N(a_{i})\cong B^{2}_{1}\times[r_{i},r_{i+1}]$ be a neighborhood of $a_{i}$ in $F\times[r_{i},r_{i+1}]$. Then, $N(a_{i})$ intersects $F_{i}$ (resp. $F_{i+1}$) in a disk $O^{-}_{i}:=B^{2}_{1}\times\{r_{i}\}$ (resp. $O^{+}_{i}:=B^{2}_{1}\times\{r_{i+1}\}$). Define $\Sigma$ as the surface obtained from $\bigcup F_{i}$ by replacing $\bigcup O^{\pm}_{i}$ with the tubes $\bigcup\partial B^{2}_{1}\times[r_{i},r_{i+1}]$. Then, $\Sigma$ cuts $P$ into two compact $3$-manifolds $V$ and $W$, each obtained from $F\times I$ and handlebodies by connecting them with $1$-handles. More precisely, if $n$ is even
$V\cong W=F\times[-1,r_{1}]\bigcup_{i~{}even}(F_{i}\setminus\mathrm{int}\,O^{-}_{i})\times[r_{i},r_{i+1}]\bigcup_{i~{}odd}N(a_{i}).$
The odd case is the same except that $V$ is a handlebody and $\partial W\cap\partial P=\partial P$. Thus, $\Sigma$ is a Heegaard surface of genus $(n+1)g$. Furthermore, as a consequence of the main theorem of Lee [28], we have
Theorem 6.1 ([28]).
The topological index of $\Sigma^{n}$ is at most $n$.
Note that when $n>1$, $\Sigma^{n}$ is stabilized and hence the index of $\Sigma^{n}$ is at least $2$. (The precise index is not known though it is likely equal to $n$.)
In this section, we see that the proof in the previous section applies to this example after slight modification: We will show
Theorem 6.2.
For every $n$, $(P,\Sigma^{n})$ satisfies the thick isotopy property.
The setting of the previous section is different from this example in the following two points: 1) $P$ has boundary; and 2) we do not know if $\pi_{d-1}(\Gamma(\Sigma))$ is finitely generated or not, where $d>0$ denotes the index of $\Sigma$.
1) is concerned with Theorem 4.1. Although the theorem assumes that the $3$-manifold is closed, the same conclusion also holds when the $3$-manifold has a boundary whose mean curvature field is inward pointing (cf. [29, Theorem 2.1]). Furthermore, we can endow $P$ with a hyperbolic metric so that $\partial P$ satisfies this condition. Indeed, $F$ can be embedded in some hyperbolic $3$-manifold as a strictly stable minimal surface, and we can identify $P$ with a small tubular neighborhood of $F$. Thus, 1) is not the issue.
As for 2), we need to show Lemma 5.3 without the assumption on $\pi_{d-1}(\Gamma(\Sigma))$. The idea of the proof is as follows: Given a sweep-out $\{\Sigma_{s}\}_{s\in B}$, we can squeeze $\{\Sigma_{s}\}_{s\in B}$ into a thin product region by an isotopy $q_{t}$ so that every slice has bounded area. Using this isotopy we can construct an extension of $\{\Sigma_{s}\}_{s\in B}$ with appropriate properties.
In what follows, we assume that $n$ is even for simplicity, although the same argument applies to an odd $n$ with straightforward modifications.
We begin with a few definitions. For $11\leq i\leq n+$ 1, put $F_{i}^{\circ}:=F_{i}\setminus\operatorname{int}O^{-}_{i}$. We say a compressing disk $D\subset F_{i}^{\circ}\times[r_{i},r_{i+1}]$ is vertical if $D=c\times[r_{i},r_{i+1}]$ for some properly embedded arc $c$ in $F_{i}^{\circ}$. Let $\{c^{1}_{j}\}^{2g}_{j=1}$ be properly embedded, pairwise disjoint arcs in $F_{1}^{\circ}$, which cut $F_{1}^{\circ}$ into a disk. Similarly, let $\{e^{2}_{j}\}^{2g}_{j=1}$ be arcs in $F_{2}^{\circ}$ satisfying the same condition. Furthermore, we may choose these arcs so that $|(c^{1}_{j}\times\{r_{2}\})\cap e^{2}_{j^{\prime}}|=\delta_{jj^{\prime}}$, as depicted in Figure 2.
[…]
Figure 2. Arcs in $F_{2}\setminus(O^{+}_{1}\cup O^{-}_{2})$: $c^{1}_{1}\times\{r_{2}\},\ldots,c^{1}_{2g}\times\{r_{2}\}$ (green) and $e^{2}_{1}\ldots,e^{2}_{2g}$ (red).
Repeating the same argument for $13\leq i\leq n+$ 1, we obtain collections of arcs $\{c^{i}_{j}\}^{2g}_{j=1}$ and $\{e^{i}_{j}\}^{2g}_{j=1}$. These arcs define vertical disks $C^{i}_{j}:=c^{i}_{j}\times[r_{i-1},r_{i}]$, $E^{i}_{j}:=e^{i}_{j}\times[r_{i},r_{i+1}]$ in $V$, $W$ respectively. By definition, $|C^{i}_{j}\cap E^{i+1}_{j^{\prime}}|=\delta_{jj^{\prime}}$. Set $\mathscr{C}:=\{C^{i}_{j}\}$ and :=$\mathscr{E}:=\{E^{i}_{j}\}$.
Suppose that $D$ and $D^{\prime}$ are (possibly isotopic) disjoint compressing disks on the same side of $\Sigma$, say $V$. Let $c\subset\partial V$ be a simple arc intersecting $D\cup D^{\prime}$ in its endpoints. We say $D^{\prime\prime}$ is obtained from $D$ and $D^{\prime}$ by a band summing along $c$ if $D^{\prime\prime}$ is obtained from $D\cup D^{\prime}$ by replacing $N(c)\cap(D\cup D^{\prime})$ with a rectangle $b=\mathrm{fr}(N(c))\setminus(D\cup D^{\prime})$. Here $N(c)$ is a neighborhood of $c$ in $V$. We call such a rectangle $b$ a band.
Lemma 6.3.
Any compressing disk $D$ in $V$ (resp. $W$) is obtained from disks in $\mathscr{C}$ (resp. $\mathscr{E}$) by a sequence of band summing.
Proof.
The proof is by induction on the intersection number between $D$ and $\mathscr{C}$. First, suppose that $D$ is disjoint from $\mathscr{C}$. Note that $\mathscr{C}$ cuts $V$ into $F\times I$ with “scars”, each of which corresponds to a foot of a $1$-handle dual to a disk in $\mathscr{C}$. So, we can view $D$ as a disk in $F\times I$ and $D$ is isotopic to a disk $D^{\prime}$ on $\partial(F\times I)$ that contains some scars. In other words, $D^{\prime}$ is a neighborhood of the union of some scars and arcs connecting them. Thus, $D$ is the result of band summing of disks in $\mathscr{C}$.
Next, suppose that $|D\cap\mathscr{C}|>0$. By the innermost disk argument, we may assume that $D$ intersects $\mathscr{C}$ only in finitely many arcs. Every arc in $D\cap\mathscr{C}$ cobounds a bigon $\Delta\subset\mathscr{C}$ together with a subarc in $\bigcup\partial\mathscr{C}$. Choose a bigon $\Delta$ such that $\mathrm{int}\,\Delta\cap D=\emptyset$. (In other words, $\Delta$ is an outermost bigon.) $\partial$-compressing $D$ along $\Delta$ yields the new compressing disks $D^{\prime}$, $D^{\prime\prime}$ which, by induction, are obtained from disks in $\mathscr{C}$ by a sequence of band summing. As $D$ is obtained by band summing from $D^{\prime}$ and $D^{\prime\prime}$, the conclusion follows. ∎
For $t\in I$, let $q_{t}:F\times[-1,1]\to F\times[-1,1]$ be the map given by $q_{t}(x,r)=(x,(1-t)r)$.
Lemma 6.4.
Suppose that $\{D_{v}\}$ is a finite collection of compressing disks for $\Sigma$. For every $\epsilon>0$, there exists $\epsilon^{\prime}>0$ satisfying the following. For every $v$, there exists a disk $D^{\prime}_{v}$ isotopic to $D_{v}$ such that for any $t\in(1-\epsilon^{\prime},1)$
$\mathrm{Area}(q_{t}(D^{\prime}_{v}))<\epsilon.$
Moreover, if $D_{v}$ and $D_{w}$ are disjoint, then the same is true for $D^{\prime}_{v}$ and $D^{\prime}_{w}$.
Proof.
Note that if $D_{v}$ is a vertical disk, then $\mathrm{Area}(q_{t}(D_{v}))\to 0$ as $t\to 1$. The idea of the proof is rather simple: By Lemma 6.3, $D_{v}$ can be expressed as a band sum of vertical disks, and thinning each band isotopes $D_{v}$ to a disk $D^{\prime}_{v}$ such that $\mathrm{Area}(q_{t}(D^{\prime}_{v}))\to 0$. We modify this argument and show that $D^{\prime}_{v}$,$D^{\prime}_{w}$ can be taken to be disjoint if $D_{v}$,$D_{w}$ are disjoint.
We start with some setups. Let #$\#C^{i}_{j}$ be a disk in $V$ obtained from $C^{1}_{j},C^{3}_{j},\ldots,C^{i}_{j}$ by band summing along arcs in $\partial E^{2}_{j},\partial E^{4}_{j},\ldots,\partial E^{i-1}_{j}$, as shown in Figure 3.
[…]
Figure 3. #$\#C^{i}_{j}$ is obtained from $C^{1}_{j},C^{3}_{j},\ldots,C^{i}_{j}$ (green) by band summing along arcs in $\partial E^{2}_{j},\partial E^{4}_{j},\ldots,\partial E^{i-1}_{j}$ (bold red).
Set $\mathscr{C}^{\#}:=\{\#C^{i}_{j}\}$. Note that $\mathscr{C}^{\#}$ is orthogonal to $\mathscr{E}$ (i.e. |#$|\#C^{i}_{j}\cap E^{i+1}_{j}|=1$ and |#$|\#C^{i}_{j}\cap E^{i^{\prime}}_{j^{\prime}}|=0$ otherwise), and from this point of view $\mathscr{C}^{\#}$ is easier to handle than $\mathscr{C}$. So, we work with $\mathscr{C}^{\#}$ in the following argument.
Observe that $C^{i}_{j}$ is recovered from #$\#C^{i}_{j}$ and #$\#C^{i-2}_{j}$ by taking band summing. By Lemma 6.3, $D_{v}$ can be written in this form:
$D_{v}=\bigcup^{m_{v}}_{p=1}F^{p}_{v}\cup\bigcup^{n_{v}}_{k=1}b^{k}_{v},$
where $F^{p}_{v}$ is a $\mathscr{C}^{\#}$- or $\mathscr{E}$-component of $D_{v}$, depending on if $D_{v}\subset V$ or $D_{v}\subset W$ (that is, $F^{p}_{v}$ is isotopic to a disk in $\mathscr{C}^{\#}$ or $\mathscr{E}$ after attaching bigons along arcs $F^{p}_{v}\cap\bigcup^{n_{v}}_{k=1}b^{k}_{v}$), and $b^{k}_{v}$ is a band. By taking the bands connecting $C^{i}_{j}$’s to be thin enough, we may assume that
(6.1) $\mathrm{Area}(q_{t}(\#C^{i}_{j}))<\frac{\epsilon}{4m_{v}}$
for all $v$ and $t$ sufficiently close to $1$. Similarly,
(6.2) $\mathrm{Area}(q_{t}(E^{i}_{j}))<\frac{\epsilon}{4m_{v}}$
for all $v$ and $t$ sufficiently close to $1$.
We define the disk $D^{\prime}_{v}$ isotopic to $D_{v}$ as follows. Let $N(\mathscr{C}^{\#})\cong\mathscr{C}^{\#}\times[-1,1]$, $N(\mathscr{E})\cong\mathscr{E}\times[-1,1]$ be small product neighborhoods of $\mathscr{C}^{\#}$ and $\mathscr{E}$, respectively. Isotope $D_{v}$ so that $F^{p}_{v}$ is a subdisk of $\mathscr{C}^{\#}\times\{u\}$ or ×$\mathscr{E}\times\{u\}$ for some $u$, depending on if $D_{v}\subset V$ or $D_{v}\subset W$. Let $F^{\prime p}_{v}$ denote the resulting disk that corresponds to $F^{p}_{v}$. If $b^{k}_{v}$ is a band of $D_{v}$, there is a unique rectangle $\overline{b}^{k}_{v}\subset\Sigma$ (the “shadow” of $b^{k}_{v}$) determined by the band sum structure of $D_{v}$.
Claim.
There exists a point $x\in\Sigma\setminus(N(\mathscr{C}^{\#})\cup N(\mathscr{E}))$ such that for every $v$ and $k$, $\overline{b}^{k}_{v}$ does not contain $x$.
Proof.
We can see this, for example, as follows. Let $S\subset\Sigma\setminus(N(\mathscr{C}^{\#})\cup N(\mathscr{E}))$ be a three-holed sphere bounded by essential simple closed curves. Suppose that $D_{v}$ intersects $S$ minimally. Then, every rectangle $\overline{b}^{k}_{v}$ intersects $S$ in finitely many pairwise disjoint rectangles. Note that there are only six possible types for such rectangles up to twisting around $\partial S$, as shown in Figure 4 (a). Consider components $\overline{b}_{1}$, $\overline{b}_{2}$ of $\overline{b}^{k}_{v}\cap S$, $\overline{b}^{l}_{w}\cap S$ respectively. If $\partial S$, $\overline{b}_{1}$ and $\overline{b}_{2}$ form a triangle, we can push it off $S$ as in Figure 4 (b). By repeating this operation until all such triangles are eliminated, we may assume that $\overline{b}_{1}$ and $\overline{b}_{2}$ do not intersect near $\partial S$ if $\overline{b}_{1}$, $\overline{b}_{2}$ are of different types. This implies $\partial S\setminus\bigcup_{k,v}\overline{b}^{k}_{v}\neq\emptyset$.
[…]
Figure 4. (a) There are only six possible types for rectangles in $S$ up to twisting around $\partial S$. (b) A triangle formed by $\partial S$ and sides of $\overline{b}_{1}$, $\overline{b}_{2}$ can be pushed off $S$.
Let $O_{x}$ be a small neighborhood $O_{x}\cong\operatorname{int}B^{2}_{1}$ of $x$ in $\Sigma$. Imagine that we expand $O_{x}$ across $\Sigma$ and push $\overline{b}^{k}_{v}$ simultaneously into a thin neighborhood of a graph $G\subset\Sigma$ depicted in Figure 5: The graph $G$ can be taken so that it satisfies the following:
  • $G$ is a deformation retract of $\Sigma\setminus O_{x}$.
  • $G$ contains the boundaries of disks #$\#C^{1}_{1},E^{2}_{1},\#C^{3}_{1},E^{4}_{1},\ldots,\#C^{n-1}_{2g},E^{n}_{2g}$ as its loops.
[…]
Figure 5. $G$ is a deformation retract of $\Sigma\setminus O_{x}$.
This map extends to an ambient isotopy that leaves $N(\mathscr{C}^{\#})$ and $N(\mathscr{E})$ invariant. Isotope further each $b^{k}_{v}$ to a rectangle $b^{\prime k}_{v}$ near $\overline{b}_{v}^{k}$ so that
(6.3) $\sum^{n_{v}}_{k=1}\mathrm{Area}(q_{t}(b^{\prime k}_{v}))<\frac{\epsilon}{2}$
for $t$ sufficiently close to $1$. Let $D^{\prime}_{v}$ be the resulting disk.
We see that $D^{\prime}_{v}$ satisfies the desired property. If $D_{v}\subset V$, by taking $F^{\prime p}_{v}$ to be close enough to #$\#C^{i}_{j}$, we may assume that
$\mathrm{Area}(q_{t}(F^{\prime p}_{v}))<\mathrm{Area}(q_{t}(\#C^{i}_{j}))+\frac{\epsilon}{4m_{v}}$
Similarly,
$\mathrm{Area}(q_{t}(F^{\prime p}_{v}))<\mathrm{Area}(q_{t}(E^{i}_{j}))+\frac{\epsilon}{4m_{v}}$
if $D_{v}\subset W$. By (6.1) or (6.2) we have
$\mathrm{Area}(q_{t}(F^{\prime p}_{v}))<\frac{\epsilon}{2m_{v}}$
for $t$ sufficiently close to $1$. Combining this with (6.3), we have
$\mathrm{Area}(q_{t}(D^{\prime}_{v}))<\sum^{m_{v}}_{p=1}\mathrm{Area}(q_{t}(F^{\prime p}_{v}))+\sum^{n_{v}}_{k=1}\mathrm{Area}(q_{t}(b^{\prime k}_{v}))<\epsilon$
for $t$ sufficiently close to $1$.
Finally, we see that if $D_{v}$, $D_{w}$ are disjoint, then so are $D^{\prime}_{v}$, $D^{\prime}_{w}$. Observe that if $F^{p}_{v}$ is already in an appropriate position, then we can isotope $D_{w}$ in the complement of $\bigcup_{p}F^{p}_{v}$ so that the same is true for $F^{q}_{w}$. The isotopies from $b^{k}_{v}$ to $b^{\prime k}_{v}$ and from $b^{l}_{w}$ to $b^{\prime l}_{w}$ can be done simultaneously. Thus, we can take $F^{\prime p}_{v}$,$F^{\prime q}_{w}$,$b^{\prime k}_{v}$ and $b^{\prime l}_{w}$ so that $F^{\prime p}_{v}\cap F^{\prime q}_{w}=b^{\prime k}_{v}\cap b^{\prime l}_{w}=\emptyset$, which proves the lemma. ∎
Let $\delta>0$ as in Section 5. Set
$\delta^{\prime}:=\min\left\{\delta,\inf_{t\in[0,1)}\gamma(q_{t}(\Sigma))\right\},$
and
$C^{\prime}:=\max_{t\in I}\mathrm{Area}(q_{t}(\Sigma))+1.$
Note that $\delta^{\prime}>0$ by Lemma 6.3.
Lemma 6.5.
Suppose that $\psi:S^{d-1}\to\Gamma(\Sigma)$ is given, and $\{\Sigma_{s}\}_{s\in B}$ is the sweep-out constructed from $\psi$ as in Step 1 of Section 5. Let $t^{\prime}\in(0,1)$. Then, there exists an extension $\{\Sigma_{ts}\}_{(t,s)\in[0,t^{\prime}]\times B}$ of $\{\Sigma_{s}\}_{s\in B}$ such that
  1. (1)
    $\Sigma_{t0}=\Sigma$ for $t\in[0,t^{\prime}]$.
  2. (2)
    $\mathrm{Area}(\Sigma_{t^{\prime}s})<C^{\prime}$ for $s\in B$.
  3. (3)
    $B\setminus U_{t^{\prime}}$ is a star-shaped region, where $U_{t^{\prime}}$ is the set of those points $s$ such that $\Sigma_{t^{\prime}s}$ is $\delta^{\prime}$-compressible.
Proof.
Fix $t\in I$ for a moment. Let $\ell_{t}:B_{1/2t}\to[0,t]$ be the map given by $\ell_{t}(s)=2|s|$. Set $B_{[1/2t,1]}=B\setminus\mathrm{int}\,B_{1/2t}$. Let $\varpi_{t}:B_{[\frac{1}{2}t,1]}\to B$ be the map given by
$\varpi_{t}(rs)=\frac{2r-t}{2-t}s$
for $r\in[1/2t,1]$ and $s\in S^{d-1}=\partial B$. Then, define the sweep-out $\{\Sigma_{ts}\}_{s\in B}$ by
$\Sigma_{ts}:=\begin{cases}q_{\ell_{t}(s)}(\Sigma)&s\in B_{\frac{1}{2}t},\\ q_{t}\left(\Sigma_{\varpi_{t}(s)}\right)&s\in B_{[\frac{1}{2}t,1]}.\end{cases}$
As $\Sigma_{ts}$ varies smoothly with $t$, we obtain the sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[0,1)\times B}$. It follows from the construction that $\Sigma_{t0}=\Sigma$ for $t\in I$.
It follows from the definition of $\delta^{\prime}$ that $B_{\frac{1}{2}t}\cap U_{t}=\emptyset$ for all $t$. On the other hand, after relabeling via $\varpi_{t}:B_{[\frac{1}{2}t,1]}\to B$, $\{\Sigma_{ts}\}_{s\in B_{[1/2t,1]}}$ is nothing but the image of $\{\Sigma_{s}\}_{s\in B}$ by $q_{t}$. Thus, the item (3) follows from Lemma 5.2.
It remains to see (2) holds. We see that for $t$ sufficiently close to $1$, $\{\Sigma_{ts}\}_{s\in B}$ can be modified so that $\{\Sigma_{ts}\}_{s\in B}$ satisfies (2).
For each vertex $v$ of $S^{d-1}$, let $D_{v}$ be a representative of $\psi(v)$ such that if $v$ and $w$ are in the same simplex, $D_{v}$ and $D_{w}$ are disjoint from each other. By Lemma 6.4, there exist $\epsilon^{\prime}>0$ and a disk $D^{\prime}_{v}$ isotopic to $D_{v}$ such that
(6.4) $\mathrm{Area}(q_{t}(D^{\prime}_{v}))<\frac{1}{4N}$
for $t\in(1-\epsilon^{\prime},1)$. Here, $N>0$ is the number of vertices of $S^{d-1}$. Fix $t\in(1-\epsilon^{\prime},1)$.
Consider the sweep-out $\Sigma^{\prime}_{s\in B}$ given as follows. For each $v$, set $D^{\prime\prime}_{v}:=q_{t}\left(D^{\prime}_{v}\right)$ and fix an identification $\iota_{v}:B^{2}_{1}\times[-1,1]\to N(D^{\prime\prime}_{v})$. By Lemma 6.4, $D^{\prime\prime}_{v}$ and $D^{\prime\prime}_{w}$ are disjoint from each other if $v$, $w$ are contained in the same simplex of $S^{d-1}$. So, we can construct a sweep-out $\{\Sigma^{\prime\prime}_{s}\}_{s\in B}$ with $\Sigma^{\prime\prime}_{0}=q_{t}(\Sigma)$ from $D^{\prime\prime}_{v}$, $N(D^{\prime\prime}_{v})$ and $\iota_{v}$ as in Step 1 of Section 5. As before, define
$\Sigma^{\prime}_{s}:=\begin{cases}q_{\ell_{t}(s)}(\Sigma)&s\in B_{\frac{1}{2}t},\\ q_{t}\left(\Sigma^{\prime\prime}_{\varpi_{t}(s)}\right)&s\in B_{[\frac{1}{2}t,1]}.\end{cases}$
By definition, $\mathrm{Area}(\Sigma^{\prime}_{s})<C^{\prime}$ for $s\in B_{1/2t}$. Thus, it suffices to see that $\mathrm{Area}(\Sigma^{\prime}_{s})<C^{\prime}$ for $s\in B_{[1/2t,1]}$.
On each simplex $\sigma$ of $B$, we can express the area of $\Sigma^{\prime}_{s}$ as
$\mathrm{Area}(\Sigma^{\prime}_{s})=\mathrm{Area}\left(\Sigma^{\prime}_{s}\setminus\bigcup_{v\in\sigma}N(D^{\prime\prime}_{v})\right)+\sum_{v\in\sigma}\mathrm{Area}\left(\Sigma^{\prime}_{s}\cap N(D^{\prime\prime}_{v})\right).$
For $\tau\in I$, let $\varsigma_{\tau}:B^{2}_{1}\times[-1,1]\to B^{2}_{1}\times[-1,1]$ be the map given by $\varsigma_{\tau}(x,u):=(x,(1-\tau)u)$. Set $\lambda^{v}_{\tau}:=\iota_{v}\circ\varsigma_{\tau}\circ\iota^{-1}_{v}$. Then, $\lambda^{v}_{\tau}$ shrinks $N(D^{\prime\prime}_{v})$ in the $I$-direction. Letting $\tau\to 1$, we have
$\mathrm{Area}(\lambda^{v}_{\tau}(\Sigma^{\prime}_{s}\cap N(D^{\prime\prime}_{v})))<2\mathrm{Area}(D^{\prime\prime}_{v})+\frac{1}{2N},$
for $s\in\sigma\cap B_{[1/2t,1]}$ and $\tau$ sufficiently close to $1$. Thus, after shrinking each $N(D^{\prime\prime}_{v})$ by $\lambda^{v}_{\tau}$, we have
$\mathrm{Area}(\Sigma^{\prime}_{s})$ $<\mathrm{Area}\left(\Sigma^{\prime}_{s}\setminus\bigcup_{v\in\sigma}N(D^{\prime\prime}_{v})\right)+\sum_{v\in\sigma}2\mathrm{Area}(D^{\prime\prime}_{v})+\sum_{v\in\sigma}\frac{1}{2N}$
$<\mathrm{Area}\left(\Sigma^{\prime}_{s}\setminus\bigcup_{v\in\sigma}N(D^{\prime\prime}_{v})\right)+1$
on $\sigma\cap B_{[1/2t,1]}$. Here, we used the inequality 6.4 for the second line. By construction, $\Sigma^{\prime}_{s}$ coincides with $q_{t}(\Sigma)$ in the complement of $\bigcup N(D^{\prime\prime}_{v})$, and hence $\mathrm{Area}(\Sigma^{\prime}_{s}\setminus\bigcup N(D^{\prime\prime}_{v}))=\mathrm{Area}(q_{t}(\Sigma)\setminus\bigcup N(D^{\prime\prime}_{v}))$. Combining this with the above inequality implies
$\mathrm{Area}(\Sigma^{\prime}_{s})<C^{\prime}$
for $s\in B_{[1/2t,1]}$.
By the same argument at the end of the proof of Lemma 5.3, we can interpolate between $\{\Sigma_{ts}\}_{s\in B}$ and $\{\Sigma^{\prime}_{s}\}_{s\in B}$ via isotopy. The resulting sweep-out satisfies (2) as well as the other two conditions. This completes the proof. ∎
Having established Lemma 6.5, Theorem 6.2 now follows by the same argument as the proof of Theorem 1.2: we have only to use Lemma 6.5 instead of Lemma 5.3. Also, a similar argument to Corollary 1.3 proves the following.
Theorem 6.6.
For every $n\in\mathbb{N}$, $\mathrm{MCG}(P,\Sigma^{n})$ is finitely generated.
Appendix A The disk complex of a Heegaard splitting of $S^{3}$
In this appendix, we prove that the disk complex of a genus $\geq 2$ Heegaard surface $\Sigma$ of $S^{3}$ is not of finite type, as mentioned in Section 1. More precisely, we show:
Proposition A.1.
Both $H_{2g-2}(\mathcal{C}(\Sigma))$ and $H_{2g-2}(\Gamma(\Sigma))$ are not finitely generated.
Here, $\mathcal{C}(\Sigma)$ is the curve complex of $\Sigma$ defined as follows. The vertices of $\mathcal{C}(\Sigma)$ are isotopy classes of essential simple closed curves on $\Sigma$. A $k$-tuple of vertices spans a $k$-simplex if these vertices admit pairwise disjoint representatives. We first prove the proposition for the curve complex, and then for the disk complex. (The assertion for the curve complex is previously known as Theorem 1.4 of Ivanov-Ji [24]. But here we reprove this fact in a way that can apply to the disk complex.)
For brevity, set $H:=H_{2g-2}(\mathcal{C}(\Sigma))$. We give the norm for $H$ as follows: For any simplicial chain $c=\sum a_{i}\sigma_{i}$, define
$\|c\|:=\sum|a_{i}|.$
For $\alpha\in H$, define
$\|\alpha\|:=\min_{[c]=\alpha}\|c\|.$
The next lemma follows from the definition.
Lemma A.2.
[…]
Take the tensor product $\mathbb{R}\otimes\widetilde{H}$. We think of $H$ as a subset of $\mathbb{R}\otimes H$.
Lemma A.3.
There exists a norm on $\mathbb{R}\otimes H$ such that
$\|r\otimes\alpha\|=|r|\|\alpha\|$
for $r\in\mathbb{R}$ and $\alpha\in H$.
Proof.
We can define, for example, the norm on $\mathbb{R}\otimes H$ as follows. (The definition is similar to the injective cross norm first introduced in [22].) Denote by $B_{H^{*}}$ the set of homomorphisms $\psi:H\to\mathbb{R}$ such that $|\psi(\alpha)|\leq\|\alpha\|$ for all $\alpha\in H$. For $\beta=\sum r_{i}\otimes\alpha_{i}\in\mathbb{R}\otimes H$, define
$\|\beta\|:=\sup\left\{\left|\sum r_{i}\psi(\alpha_{i})\right|\mid\psi\in B_{H^{*}}\right\}.$
It follows from the definition that this is a seminorm and satisfies $\|r\otimes\alpha\|=|r|\|\alpha\|$ for $r\in\mathbb{R}$ and $\alpha\in H$. If $\beta=\sum r_{i}\otimes\alpha_{i}$ and $\|\beta\|=0$, then
$|r_{i}|\|\alpha_{i}\|=\left|r_{i}\psi_{\alpha_{i}}(\alpha_{i})\right|\leq\left\|\sum r_{i}\otimes\alpha_{i}\right\|=0$
for all $i$. Here, $\psi_{\alpha_{i}}\in B_{H^{*}}$ is the homomorphism given by $\psi_{\alpha_{i}}(n\alpha_{i})=n\|\alpha_{i}\|$ for $n\in\mathbb{Z}$ and $\psi_{\alpha_{i}}=0$ otherwise. Thus, $\beta=0$ and we conclude that this seminorm is in fact a norm. ∎
Now we prove that $H$ is not finitely generated. Denote by $\mathcal{B}(r_{0})\subset\mathbb{R}\otimes H$ the ball of diameter at most $r_{0}>0$. Note that $H\neq 0$ by [24, Theorem 1.3] (or by [5, Theorem 1.1]). By taking $r_{0}$ to be large enough, we may assume that $H\cap\mathcal{B}(r_{0})$ contains a nontrivial homology class. Suppose, contrary to our claim, that $H$ is finitely generated. Then, $\mathbb{R}\otimes H$ is also finitely generated, and hence $\mathbb{R}\otimes H$ is equivalent to $\mathbb{R}^{n}$ with the Euclidean norm. It follows that $H\cap\mathcal{B}(r_{0})$ is a finite set since any closed bounded set in $\mathbb{R}\otimes H\cong\mathbb{R}^{n}$ is compact.
It follows from Lemma A.2 that $\mathrm{MCG}(\Sigma)$ maps $H\cap\mathcal{B}(r_{0})$ to itself. Thus, for every point $\alpha\in H\cap\mathcal{B}(r_{0})$, the stabilizer subgroup of $\alpha$ is an infinite group. But Corollary 5.3 in Irmer [23] says that every stabilizer subgroup for the action of $\mathrm{MCG}(\Sigma)$ on $H\setminus\{0\}$ must be trivial or $\mathbb{Z}/2\mathbb{Z}$ (the latter case occurs only when $g=2$), a contradiction.
Next, we see that $H_{2g-2}(\Gamma(\Sigma))$ is not finitely generated. The proof is similar to the curve complex case. Note that the disk complex $\Gamma(\Sigma)$ can be identified with a subcomplex of $\mathcal{C}(\Sigma)$ in the following sense. Consider the map given by sending $[D]\in\Gamma(\Sigma)$ to $[\partial D]\in\mathcal{C}(\Sigma)$. (This is not an injection because $\partial D$ may bound a disk opposite to $D$.) This map is a homotopy equivalence between $\Gamma(\Sigma)$ and its image $\mathcal{D}_{\Sigma}$. Let $H_{\mathcal{D}}$ be the subspace of $H$ consisting of those homology classes $\alpha$ such that $\alpha$ is represented by a cycle in $\mathcal{D}_{\Sigma}$.
In [5], Broaddus identified a generator of $H$ as a $\mathbb{Z}\mathrm{MCG}(\Sigma)$-module, which is represented by a nontrivial $(2g-2)$-sphere in $\mathcal{C}(\Sigma)$. Surprisingly, such a sphere can be found within $\mathcal{D}_{\Sigma}$. See e.g. Figures 1-3 in [7]. Thus, $H_{\mathcal{D}}\neq 0$. In particular, $H_{\mathcal{D}}\cap\mathcal{B}(r_{0})$ contains a nontrivial homology class for $r_{0}$ sufficiently large.
As before, suppose that $H_{2g-2}(\Gamma(\Sigma))$ is finitely generated. Then, so is $H_{\mathcal{D}}$ since $H_{\mathcal{D}}$ can be thought of as a subspace of $H_{2g-2}(\mathcal{D}_{\Sigma})\cong H_{2g-2}(\Gamma(\Sigma))$. By the same argument as above, it follows that $H_{\mathcal{D}}\cap\mathcal{B}(r_{0})$ is a finite set. Note that restricting maps $(S^{3},\Sigma)\to(S^{3},\Sigma)$ on $\Sigma$ defines the injection $\mathrm{MCG}(S^{3},\Sigma)\to\mathrm{MCG}(\Sigma)$. So we can think of $\mathrm{MCG}(S^{3},\Sigma)$ as a subgroup of $\mathrm{MCG}(\Sigma)$. By Lemma A.2, $\mathrm{MCG}(S^{3},\Sigma)$ maps $H_{\mathcal{D}}\cap\mathcal{B}(r_{0})$ to itself. As $\mathrm{MCG}(S^{3},\Sigma)$ is an infinite group, the stabilizer subgroup of a point in $H_{\mathcal{D}}\cap\mathcal{B}(r_{0})$ is an infinite group, again contradicting Corollary 5.3 in [23]. This completes the proof of Proposition A.1. ∎
As a consequence of Proposition A.1 and the main theorem of Appel [2] or Campisi-Torres [7], we have
Theorem A.4.
Suppose that $\Sigma$ is a Heegaard surface for $S^{3}$ of genus $g\geq 2$. Then the disk complex $\Gamma(\Sigma)$ is homotopy equivalent to the bouquet of countably infinitely many spheres of dimension $22g-$ 2.